OpenStudy (anonymous):

What is the standard form of a quadratic equation? What is the standard form of a quadratic equation? @Mathematics

OpenStudy (anonymous):

ax^2 + bx + c = 0, where a, b, c are constants, a is non-zero

OpenStudy (anonymous):

ok thanks. I have to make up a equation for my class would -3x^2+30x-15=0 be in standard form? I cant never made up one and it work out right.

OpenStudy (anonymous):

yes, it certainly would be.

OpenStudy (anonymous):

one way to think of one which "works out right" is to come up with the answers then work backwards to form the equation x = 2, x = 3: (x - 2)(x - 3) = 0 then expand the left hand side: x^2 - 5x + 6 = 0

OpenStudy (anonymous):

lol i just threw some numbers together. this is my first day on learning this stuff. does it matter what numbers i use>